Reports 1-1 of 1 Clear search Modify search
MIF (Noise Budget)
takafumi.ushiba - 19:47 Wednesday 19 November 2025 (35601) Print this report
Template preparationfor VIS noise budget

I started to make diaggui template for VIS noise budget.
Template file is stored at /users/Commissioning/data/NoiseBudget/TFs/template/.

For local dampling control noise, I set the excitation amlplitude so that the excitation signals are roughly 10 times larger than sensor noise aboe 10 Hz.
For actuator noise, I set the excitation amplitude so that the COILOUTF output spectrum is about 1 cnt/rtHz above 10 Hz.

I will continue the preparation work tomorrow.

Comments to this report:
takafumi.ushiba - 22:21 Thursday 20 November 2025 (35610) Print this report

I finished to make the diaggui template for ETMX.
I changed the amplitude for actuator noise projection as follows:
1. For the actuator with 0- or 1-stage of dewhitening filters, amplitude is set as 1.
2. For the actuator with 2- or 3-stage of dewhitening filters, amplitude is set as 0.1.

Since amplitude of 1 and 0.1 is corresponding to 3e-2 and 3e-3 cnt/rtHz at COILOUTF input (9e-6 and 9e-7 V/rtHz at coil driver output if there is no gain at COILOUTF), it should be at least 10 times larger than DAC noise and coil driver noise.

After amplitude tuning, I started to run the diaggui files with the automeasurement script developed by Yamamoto-kun.
To avoid turning on the OBS_INTENT after the measurement, I modified the script before runing.
I will check the results tomorrow.

Note:

If the measurement won't be finished by tomorrow morning, the request state for LSC_LOCK will be WORKING_WITH_OBSERVATION_WITHOUT_LINES (if the measurement will be finished, OBSERVATION should be requested automatically).
So, if the state won't be OBSERVATION tomorrow morning, please kill the process on k1ctr2 before starting the work.

shinji.miyoki - 1:22 Friday 21 November 2025 (35613) Print this report

When I checked IFO around 22:00, IFO was unlocked, and "WORKING_WITH_OBSERVATION_WITHOUT_LINES" was selected in the LSC_LOCK Guardian. So, the measurement seemed to fail according to this klog post  explanation "If the measurement won't be finished by tomorrow morning, the request state for LSC_LOCK will be WORKING_WITH_OBSERVATION_WITHOUT_LINES ".

So, I requested PRFPMI_RF_LOCK, and found an ugly AS image. So I set this state for several decades of minutes, and try  "WORKING_WITH_OBSERVATION_WITHOUT_LINES". However, IFO was unlocked around OMC treatment, the final stage of the "INCREASE_LAS_POWER", etc.

By the way, at the final stage of "INCREASE_LAS_POWER", the value of the laser power was not set at its maximum, such as 15.6W, but fluctuated between 15.2 and 15.6 for a while. Is this healthy situation or not? Actually, during this fluctuation, IFO was unlocked.

After this unlock, GRX trans shape became not TEM00 and its value became ~ 0.6. So some mirror alignment can be changed. (Fig.1)

Images attached to this comment
shinji.miyoki - 1:42 Friday 21 November 2025 (35614) Print this report

I touched PR3 and ETMX OpticAlign. However, I recovered the values for the P/Y of PR3 because there was no changes in GRX trans power, while I found ETMX Y should be from 1.9 to -1.0 to obtain ~1 GRX trans Norm.

If these SDF should be accepted, please accept. Of course, LSC_LOCK guardian restarted from the value of 1.9 for ETMX Y after the IFO down. However, this condition of 1.9 for ETMX Y never succeeded in reaching PRFPMI_RF_LOCK.

Anyway, after this ETMX Y value change from 1.9 to -1.0, I tried PRFPMI_RF_LOCk and it succeeded after two times failure. 

After that, I requested "WORKING_WITH_OBSERVATION_WITHOUT_LINE", and finally IFO reached this state. 

takafumi.ushiba - 13:26 Friday 21 November 2025 (35621) Print this report

I checked the measured data and several TF seems to have significantly high coherence at several frequency bands as shown in the floowing.
The other TFs are not well measured due to the lockloss or have low coherence at 10-100 Hz band.

To confirm these noises really limits the current sensitivity, noise projection using these TFs are necessary.

Fig1 (BF_COILOUTF_GAS):

This measurement is estimating BF GAS coil driver (DAC) noise, so the coil driver noise of BF GAS might affect the sensitivity around 20 and 40 Hz.

Fig2 (GASMODAL_DAMP_M2):

This measurement is estimating the feedback noise of GASMODAM_DAMP_M2 filter, so this loop might affect the current sensitivity around 20 Hz.

Fig3 (MN_COILOUTF_H3):

This measurement is estimating MN H3 coil driver (DAC) noise, so the coil driver noise of MN H3 might affect the sensitivity around 30, 40, 50 Hz.

Fig4 (MN_COILOUTF_V1):

This measurement is estimating MN V1 coil driver (DAC) noise, so the coil driver noise of MN V1 might affect the sensitivity around 30-60 Hz.

Fig5 (MN_COILOUTF_V2):

This measurement is estimating MN V2 coil driver (DAC) noise, so the coil driver noise of MN V2 might affect the sensitivity around 40 and 50 Hz.

Fig6 (MN_COILOUTF_V3):

This measurement is estimating MN V3 coil driver (DAC) noise, so the coil driver noise of MN V3 might affect the sensitivity around 30-60 Hz.

Fig7 (MN_DAMP_L):

This measurement is estimating the feedback noise of MN_DAMP_L filter, so this loop might affect the current sensitivity around below 40 Hz.

Fig8 (MN_DAMP_T):

This measurement is estimating the feedback noise of MN_DAMP_T filter, so this loop might affect the current sensitivity around below 40 and around 50 Hz.

Fig9 (MN_MNOLDAMP_P):

This measurement is estimating the feedback noise of MN_MNOLDAMP_P filter, so this loop might affect the current sensitivity around below 20, 30-40, and around 50 Hz.

Fig10 (MN_MNOLDAMP_R):

This measurement is estimating the feedback noise of MN_MNOLDAMP_R filter, so this loop might affect the current sensitivity around 50 and 65 Hz.

Fig11 (MN_MNOLDAMP_Y):

This measurement is estimating the feedback noise of MN_MNOLDAMP_Y filter, so this loop might affect the current sensitivity below 40 Hz.

Images attached to this comment
Search Help
×

Warning

×